OpenStudy (anonymous):

Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution. Equation 1 8 - 3x = 5x + 16 Equation 2 9x - 14 = 8x - 15 The solution to equation #1 is larger than the solution to equation #2. The solution to equation #1 is smaller than the solution to equation #2. The solution to equation #1 is the same as the solution to equation #2. None of the statements above describe solutions to equations shown.

OpenStudy (anonymous):

your job is so solve both can you do it?

OpenStudy (anonymous):

I think so.

OpenStudy (anonymous):

then you can answer the question asked

OpenStudy (anonymous):

Okay I need help with the first step on the first problem.

OpenStudy (anonymous):

I always get stuck on the first step.

OpenStudy (anonymous):

\[8 - 3x = 5x + 16 \] you want all the \(x\)s on one side. lets put them on the right by adding \(3x\) to both sides

OpenStudy (anonymous):

you get \[8-3x+3x=5x+3x+16\] or with practice you go right to \[8=8x+16\] then all the \(x\)s are on the right (there are 8 of them) so you want the numbers on the left. subtract \(16\) from both sides

OpenStudy (anonymous):

you get \[8-16=8x\] or \[-8=8x\] and then your last step is to divide both sides by \(8\)

OpenStudy (anonymous):

So it would be 1?

OpenStudy (anonymous):

actually it would be \(-1\) because \(-8\div 8=-1\)

OpenStudy (anonymous):

Oh.

OpenStudy (anonymous):

Okay so now 9x-14=8x-15 .. oh boy

OpenStudy (anonymous):

We would add 8x to both sides?

OpenStudy (anonymous):

careful here

OpenStudy (anonymous):

you have \(8x\) on the right, so to get rid of it you want to subtract \(8x\) from both sides, not add

OpenStudy (anonymous):

1x-14=-15?

OpenStudy (anonymous):

the reason we added in the first one is because you had \(-3x\) so to get rid of that we added because \(-3x+3x=0\) but this time we want \(8x-8x=0\) so you need to subtract

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

now the \(x\) is on the left, you want the numbers on the right

OpenStudy (anonymous):

:D whoo. Okay so now.. we add 15 or subtract from both sides?

OpenStudy (anonymous):

on the left is \(-14\) so to get rid of that you will ...

OpenStudy (anonymous):

Oh I see we are trying to get x by it's self.

OpenStudy (anonymous):

hold on make sure the plan is clear. you want the variable on one side, the numbers on the OTHER side

OpenStudy (anonymous):

yes, exactly

OpenStudy (anonymous):

So we will minus -14 on both sides?

OpenStudy (anonymous):

No wait. Add.

OpenStudy (anonymous):

then it would be 1x=1

OpenStudy (anonymous):

Then divide 1 by one and get 1

OpenStudy (anonymous):

yes add that is right but be careful of the arithmetic

OpenStudy (anonymous):

\(-15+14=...\)

OpenStudy (anonymous):

Aw darn. so it's -1

OpenStudy (anonymous):

right

OpenStudy (anonymous):

and the answer is -1

OpenStudy (anonymous):

and one more thing, if you want to make sure your teacher thinks you are paying attention, never write \(1x\) just write \(x\)

OpenStudy (anonymous):

So our answer would be the 3rd statement both equations are the same.

OpenStudy (anonymous):

so you do not have to divide by 1, which is like doing nothing in any case you get \(x=-1\) not \(1x=-1\) although of course they are the same

OpenStudy (anonymous):

yes, the third one i had to scroll way the hell up to see it so it took me a minute

OpenStudy (anonymous):

I see I see (:

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

you will get this, do not fret